OpenStudy (anonymous):

I need the answer, Solve for x: −4|x + 5| = −16

OpenStudy (lynfran):

lets work it out together ok

OpenStudy (anonymous):

okay

OpenStudy (lynfran):

\[-4*\left| x+5 \right|=-16\] the absolute value of x+5 is x+5, so,\[-4(x+5)=-16\]...multiply out the bracket we get\[-4x-20=-16\].. can u finish from here?

OpenStudy (lynfran):

-4x=-16+20 then divide by -4 to get x=??

OpenStudy (anonymous):

okay i got positive 9 @LynFran

OpenStudy (lynfran):

-4x=4 divide by -4 x=-1

OpenStudy (anonymous):

so its x=-1, x=9